# Bayesian Inference General Procedures for A Single-subject Test Study

### v4

@@ Abstract
Abnormality detection in identifying a single-subject which deviates from the majority of a control group dataset is a fundamental problem. Typically, the control group is characterised using standard Normal statistics, and the detection of a single abnormal subject is in that context. However, in many situations, the control group cannot be described by Normal statistics, making standard statistical methods inappropriate. This paper presents a Bayesian Inference General Procedures for A Single-Subject Test (BIGPAST) designed to mitigate the effects of skewness under the assumption that the dataset of the control group comes from the skewed Student $t$ distribution. BIGPAST operates under the null hypothesis that the single-subject follows the same distribution as the control group. We assess BIGPAST’s performance against other methods through simulation studies. The results demonstrate that BIGPAST is robust against deviations from normality and outperforms the existing approaches in accuracy. BIGPAST can reduce model misspecification errors under the skewed Student $t$ assumption. We apply BIGPAST to a Magnetoencephalography (MEG) dataset consisting of an individual with mild traumatic brain injury and an age and gender-matched control group, demonstrating its effectiveness in detecting abnormalities in a single-subject.
Keywords— Bayesian inference, Skewed Student’s $t$ distribution, Single-subject test, Magnetoencephalography (MEG), Jeffery’s prior, and False discovery rate
@@ 1 Introduction
Abnormality detection is a fundamental task in many scientific fields, including medicine, psychology, and neuroscience (Scholz and Stephens, 1987; Crawford and Howell, 1998; Crawford and Garthwaite, 2007). One example of such abnormality detection lies in the detection of traumatic brain injury (TBI), which is a fundamental problem, with over 2.8 million individuals a year in America presenting at an Emergency room (Taylor et al., 2017). We exemplify the use of skewed Student $t$ statistics in the study of brain activity as a potential biomarker for the identification of a brain injury. We use magnetoencephalography (MEG) to measure brain activity and compare single individual to normal controls from age and gender-matched cohorts. Existing literature typically assumes that the dataset from the control group follows a Normal distribution. However, extensive practical evidence shows that the Normal distribution assumption is often violated. This violation is frequently observed in brain Magnetoencephalography (MEG) datasets. Therefore, the current approaches using Normal statistics could not be directly applied. Unlike the existing literature, we assume that the dataset from the control group belongs to a more general distributional family - the skewed Student $t$ distribution, which includes the Normal distribution and Student $t$ as special cases. The benefit of this skewed Student $t$ distributional assumption is that it could cover a variety of empirical behaviours and is thus more suitable for practical needs. This generalisation is essential since many brain Magnetoencephalography (MEG) datasets show a significant level of asymmetry in the observation distribution and hence are highly skewed. It is also the case that the high skewness not only appears in a MEG dataset but also in other fields, such as cybersecurity (Kim and Song, 2023), medicine (Lu and Huang, 2017), and neuroscience (Buzsáki and Mizuseki, 2014) to name only a few.
Under the assumption of the skewed Student $t$ distribution in this article, we describe the statistical issues involved in the comparison of a single individual with a large control group. The statistical problems arise from the interplay between the mechanism of data observation from a single-subject, the type of hypothesis (two-sided, less, greater), and the sign of the skewness parameter, which we will discuss one by one in the following paragraphs.
Under the assumption that the dataset observed from the control group follows a Normal distribution, several methods exist to compare a single-subject against a control group. Crawford and Howell (1998) proposed a $t$-score test defined as follows:
@@
In this paper, we consider three kinds of alternative hypotheses: “less than” ($<$), “greater than” ($>$) and “two-sided”. For clarity, we refer to the alternative hypothesis with symbol ($<$) as the one-sided test (less), and the alternative hypothesis with symbol ($>$) as the one-sided test (greater). Under the assumption of symmetric distributions (e.g., Normal or Student $t$), one-sided tests, either “less” or “greater”, share the same absolute critical value given the same significance level. However, under the assumption of the skewed Student $t$ distribution, the absolute critical value of the one-sided test (less) is different from the absolute critical value of the one-sided test (greater) due to the relationship between skewness and heavy tails. Hence, under the assumption of a skewed Student $t$ distribution, there are six combinations of hypothesis tests: the one-sided test (less), the one-sided test (greater), and the two-sided test, each with either a positive or negative skewness parameter.
@@
Let us look into the parameter settings in Crawford et al. (2006). They set $\alpha<0$ and only considered the one-sided test (less). The observations of the control group and single-subject are generated from the same skewed Student $t$ distribution with the same skewness parameter. Furthermore, the observations of the single-subject are all negative, as described in Table 1. They used the $t$ statistic to test the abnormalities of the single-subject. Figure 1(b) describes a similar skewness parameter setting as in Crawford et al. (2006). Given the significance level 0.05, the critical value of the one-sided test (less) for the $t$ test is $t_{0.05}=-2.015$ while the critical value of the one-sided test (less) for the skewed $t$ test is $t_{0.05}^{*}=-2.568$. Apparently, $t_{0.05}>t^{*}_{0.05}$, which implies that if some single-subject observation $x^{*}$ is claimed as abnormal by $t_{0.05}^{*}$, then $x^{*}$ must be claimed as abnormal by $t_{0.05}$.
@@
Let’s define the model misspecification error for the one-sided test (less) as the probability $\Pr_{t}(t_{0.05}^{*}\leq x\leq t_{0.05})$ where $\Pr_{t}(\cdot)$ is the probability under the Student $t$ distribution. In Figure 1(b), the model misspecification error for the one-sided test (less) is $\Pr_{t}(-2.568\leq x\leq-2.015)\approx 0.0249$. However, if we consider the one-sided test (greater) under the same settings, the model misspecification error for the one-sided test (greater) is $\Pr_{t}(0.0370\leq x\leq 2.015)\approx 0.3131$. The model misspecification error for the one-sided test (greater) is almost 12 times higher than that for the one-sided test (less). This means that the one-sided test (less) is more robust than the one-sided test (greater) when $\alpha<0$. Under the settings of $\alpha<0$, one-sided test (less) and not well-mixed observations of single-subject, Crawford et al. (2006) claimed that the $t$ score is robust in terms of Type I error even under skewed Student $t$ distribution. However, this assertion does not hold if the one-sided test (less) is replaced with the one-sided test (greater). For the power comparison, Crawford and Garthwaite (2006) make a similar statement. However, by using a similar discussion to the above, we can find that the robustness statement for the power only holds for the specific settings in their paper.
@@
We can conclude that the interplay between skewness (whether positive or negative), the type of hypothesis (‘two-sided’, ‘less’, or ‘greater’), and the single-subject observations (whether well-mixed or not) significantly influences both the Type I error rate and the power of the $t$ score test under the assumption of the skewed Student $t$ distribution. The key contribution is that this paper introduces a Bayesian Inference General Procedures for A Single-subject Test (BIGPAST) framework, which employs the nested sampling technique under the skewed Student $t$ assumption to fully discuss the test performance in different settings (i.e., the combinations of the sign of $\alpha$, type of hypothesis and well-mixed single-subject observations).
Additionally, another contribution of this paper lies in the fact that we propose a Jeffreys prior, which is defined by explicit mathematical formulas. We evaluate BIGPAST’s performance against the existing approaches through simulation studies. The results demonstrate that BIGPAST is robust against departures from normality and surpasses the existing methods based on the assumption of a Normal distribution in terms of Type I error, power and accuracy.
@@ 19       $i\leftarrow i+1$;
In Algorithm 1, the Metropolis-Hastings algorithm draws samples from the posterior distribution. The step size $\delta$ is set to 0.5, and the burn-in rate $b$ is set to 0.4. The algorithm returns the samples after the burn-in period. $u_{\nu},u_{\omega},u$ are random numbers drawn from the uniform distribution $U(0,1)$. $\epsilon_{\alpha}$ and $\epsilon_{\xi}$ are drawn from the Normal distribution with mean zero and standard deviation $\delta$. As $\nu$ and $\omega$ are positive, the proposal distributions for $\nu$ and $\omega$ are the truncated normal distributions with mean $\nu^{\text{old}}$, $\omega^{\text{old}}$ and standard deviation $\delta$, respectively, see lines 6-9 of Algorithm 1. The proposal distributions for $\alpha$ and $\xi$ are normal distributions with mean $\alpha^{\text{old}}$, $\xi^{\text{old}}$ and standard deviation $\delta$, respectively, see lines 10-11 of Algorithm 1. The truncated proposal distributions do not have symmetric densities, so we have to adjust the acceptance rate $A$ by adding the logarithm of the ratios, i.e., $p_{\nu}$ and $p_{\omega}$ to the acceptance rate $A$, see lines 12-14 of Algorithm 1.
@@ 2.2 Jeffreys Priors
In the literature, two types of priors are commonly used for $\xi$ and $\omega$ in Equation 2: the partial information prior (Sun and Berger, 1998) and the independence Jeffreys prior (Rubio and Steel, 2014). Both of these priors assume that the joint prior of $\xi$ and $\omega$, denoted as $\boldsymbol{\pi}(\xi,\omega)$, is proportional to $\omega^{-1}$. Consequently, the joint prior of $\alpha,\nu,\xi,\omega$ can be expressed as follows:
@@ 3.3 Model Misspecification Error
In this section, we conduct a comparative analysis of the BIGPAST, CG, and AD methodologies under varying alternative hypotheses and parameter configurations. The predetermined significance level is 0.05. The skewness parameter $\alpha$ and degrees of freedom $\nu$ are assigned values from the sets $(1,10)$, $(1,5)$, $(2,5)$, and $(3,5)$, respectively. The location and scale of the skewed Student $t$ distribution are set to 0 and 1, respectively. Other sample settings include a sample size $n=100$, $c_{\nu}=1$, and a Metropolis-Hastings sampling size of 2000. The comparison procedures are as follows: Firstly, $N=200$ independent samples are randomly generated from the skewed Student $t$ distribution with parameters $\alpha$ and $\nu$. Secondly, $m=1000$ single-subjects are randomly drawn, and the BIGPAST, CG, and AD tests are conducted on each of the $N$ samples. Thirdly, for each of $N$ samples, the Type I error (false positive rate), power (true positive rate), and accuracy are calculated based on the $m=1000$ single-subjects. Finally, the Type I error, power, and accuracy are reported in the form of boxplots for the four skewed Student $t$ parameter settings, as shown in Figures 11, 12, 13 and 2.
@@
We utilise the total variation distance to quantify the disparity between skewed Student $t$ (for BIGPAST) and Normal (for CG) distributions. The total variation distances of the aforementioned four settings are: 0.057, 0.111, 0.155 and 0.186, respectively, as illustrated in Figures 11(a), 12(a), 13(a) and 2(a). In the context of a two-sided hypothesis test, the divergence between the BIGPAST and CG methodologies becomes larger as the total variation distance increases, as shown in Figures 11, 12, 13 and 2. The result of the AD approach demonstrates robustness for different total variation distances but performs significantly worse than BIGPAST in terms of Type I error, as seen in Figure 2(b). The AD approach also performs slightly worse than BIGPAST in terms of accuracy, as depicted in Figure 2(d).
@@
The complexity of the one-sided test (either ‘greater’ or ‘less’) compared to the two-sided test due to the skewed Student $t$ distribution increases due to skewed Student $t$ distribution having light and heavy tails. The model misspecification error impacts the tails of the skewed Student $t$ distribution differently. When $\alpha>0$ (resp. $\alpha<0$), the skewed Student $t$ distribution has a right (resp. left) heavy tail. In this paragraph, we explore the scenario where the direction of the alternative hypothesis is ‘greater’, i.e., one-sided test (greater). If we implement the CG test, the 95% quantile of Normal distribution is less than the 95% quantile of skewed Student $t$ distribution. Consequently, the Type I error of the CG test exceeds that of the BIGPAST test, leading to a higher power for the CG test, as shown in Figures 2(b) and 2(c). This can be explained by the light-red region between the 95% quantile of skewed Student $t$ distribution and the 95% quantile of Normal distribution, as depicted in Figure 4(b). Theoretically, all single-case observations from this region will be classified as positive by the CG test and as negative by the BIGPAST test, leading us to refer to this region as the false positive region. The error generated by the false positive region arises from model misspecification rather than sampling randomness, as illustrated in Figure 4.
@@ 4 Real Data Analysis
This dataset is derived from a mild traumatic brain injury (mTBI) study conducted by Innovision IP Ltd. It comprises a functional time series of Magnetoencephalography (MEG) data and an anatomical T1-weighted Magnetic Resonance Imaging (MRI) scan for 103 healthy subjects and one patient with mild traumatic brain injury. The MEG data, recorded from 306 sensors, was pre-processed using the ‘MNE-Python’ package (Gramfort et al., 2013). The MEG data was analysed to reconstruct the source generators using a minimum norm inverse method. The MRI data was pre-processed with the ‘FreeSurfer’ package (Han et al., 2006). The standard FreeSurfer pipeline (detailed on their website) was followed and includes motion correction, intensity normalisation, skull-stripping, normalisation to template space and segmentation of the cortex and subcortical structures using the Desikan-Killiany atlas. This results in 68 regions covering the entire cortex. For each cortical region, we calculate the average of source powers within this region. This process yields a control group dataset with dimensions $103\times 68\times 100$ and a patient dataset with dimensions $68\times 100$. To simplify, we treat the epochs as repeated observations and take averages over them to reduce the data’s complexity. Consequently, the control group dataset has dimensions $103\times 68$, and the patient dataset is reduced to a vector of length $68$.

### v5

@@ Abstract
Abnormality detection in identifying a single-subject which deviates from the majority of a control group dataset is a fundamental problem. Typically, the control group is characterised using standard Normal statistics, and the detection of a single abnormal subject is in that context. However, in many situations, the control group cannot be described by Normal statistics, making standard statistical methods inappropriate. This paper presents a Bayesian Inference General Procedures for A Single-subject Test (BIGPAST) designed to mitigate the effects of skewness under the assumption that the dataset of the control group comes from the skewed Student $t$ distribution. BIGPAST operates under the null hypothesis that the single-subject follows the same distribution as the control group. We assess BIGPAST’s performance against other methods through simulation studies. The results demonstrate that BIGPAST is robust against deviations from normality and outperforms the existing approaches in accuracy nearest to the nominal accuracy 0.95. BIGPAST can reduce model misspecification errors under the skewed Student $t$ assumption by up to 12 times, as demonstrated in Section 3.3. We apply BIGPAST to a Magnetoencephalography (MEG) dataset consisting of an individual with mild traumatic brain injury and an age and gender-matched control group. For example, the previous method failed to detect abnormalities in 8 brain areas, whereas BIGPAST successfully identified them, demonstrating its effectiveness in detecting abnormalities in a single-subject.
Keywords— Bayesian inference, Skewed Student’s $t$ distribution, Single-subject test, Magnetoencephalography (MEG), Jeffreys prior
@@ 1 Introduction
Abnormality detection is a fundamental task in many scientific fields, including Medicine, Psychology, Econometrics, Telehealthcare, Neuroscience as well as other fields for Industry 5.0 (Crawford and Garthwaite, 2007; Das et al., 2023; Higbee and McDonald, 2024; Kim and Song, 2023; Ktena et al., 2024; Yin et al., 2024). One example of such abnormality detection lies in the detection of traumatic brain injury (TBI), which is a fundamental problem, with over 2.8 million individuals a year in America presenting at an Emergency room (Taylor et al., 2017). We exemplify the use of skewed Student $t$ statistics in the study of brain activity as a potential biomarker for the identification of a brain injury. We use magnetoencephalography (MEG) to measure brain activity and compare single individual to normal controls from age and gender-matched cohorts. Existing literature typically assumes that the dataset from the control group follows a Normal distribution. However, extensive practical evidence shows that the Normal distribution assumption is often violated. This violation is frequently observed in brain Magnetoencephalography (MEG) datasets. Therefore, the current approaches using Normal statistics could not be directly applied. Unlike the existing literature, we assume that the dataset from the control group belongs to a more general distributional family - the skewed Student $t$ distribution, which includes the Normal distribution and Student $t$ as special cases. The benefit of this skewed Student $t$ distributional assumption is that it could cover a variety of empirical behaviours and is thus more suitable for practical needs. This generalisation is essential since many brain Magnetoencephalography (MEG) datasets show a significant level of asymmetry in the observation distribution and hence are highly skewed. It is also the case that the high skewness not only appears in a MEG dataset but also in other fields, such as Cybersecurity (Kim and Song, 2023), Medicine (Ktena et al., 2024), Econometric (Higbee and McDonald, 2024), Neuroscience (Zimmermann et al., 2024), Telehealthcare (Das et al., 2023) and Internet of Thing (IoT) (Yin et al., 2024) to name only a few.
The main focus of this research is the adaptability of statistical models under the skewed student $t$ distribution. Specifically, the statistical challenges arise from the interplay between the mechanism of data observation from a single-subject, the type of hypothesis (two-sided, less, greater), and the sign of the skewness parameter, which we will discuss one by one in the following paragraphs.
The mechanism of data observation in Type I error and Power test. Under the assumption that the dataset observed from the control group follows a Normal distribution, several methods exist to compare a single-subject against a control group. Crawford and Howell (1998) proposed a $t$-score test defined as follows:
@@
The effect of hypothesis (two-sided, less, greater). In this paper, we consider three kinds of alternative hypotheses: “less than” ($<$), “greater than” ($>$) and “two-sided”. For clarity, we refer to the alternative hypothesis with symbol ($<$) as the one-sided test (less), and the alternative hypothesis with symbol ($>$) as the one-sided test (greater). Under the assumption of symmetric distributions (e.g., Normal or Student $t$), one-sided tests, either “less” or “greater”, share the same absolute critical value given the same significance level. However, under the assumption of the skewed Student $t$ distribution, the absolute critical value of the one-sided test (less) is different from the absolute critical value of the one-sided test (greater) due to the relationship between skewness and heavy tails. Hence, under the assumption of a skewed Student $t$ distribution, there are six combinations of hypothesis tests: the one-sided test (less), the one-sided test (greater), and the two-sided test, each with either a positive or negative skewness parameter.
@@
The interaction effect of hypothesis type and skewness sign. Let us look into the parameter settings in Crawford et al. (2006). They set $\alpha<0$ and only considered the one-sided test (less). The observations of the control group and single-subject are generated from the same skewed Student $t$ distribution with the same skewness parameter. Furthermore, the observations of the single-subject are all negative, as described in Table 1. They used the $t$ statistic to test the abnormalities of the single-subject. Figure 1(b) describes a similar skewness parameter setting as in Crawford et al. (2006). Given the significance level 0.05, the critical value of the one-sided test (less) for the $t$ test is $t_{0.05}=-2.015$ while the critical value of the one-sided test (less) for the skewed $t$ test is $t_{0.05}^{*}=-2.568$. Apparently, $t_{0.05}>t^{*}_{0.05}$, which implies that if some single-subject observation $x^{*}$ is claimed as abnormal by $t_{0.05}^{*}$, then $x^{*}$ must be claimed as abnormal by $t_{0.05}$.
@@
Motivation to reduce the model misspecification error. Let’s define the model misspecification error for the one-sided test (less) as the probability $\Pr_{t}(t_{0.05}^{*}\leq x\leq t_{0.05})$ where $\Pr_{t}(\cdot)$ is the probability under the Student $t$ distribution. In Figure 1(b), the model misspecification error for the one-sided test (less) is $\Pr_{t}(-2.568\leq x\leq-2.015)\approx 0.0249$. However, if we consider the one-sided test (greater) under the same settings, the model misspecification error for the one-sided test (greater) is $\Pr_{t}(0.0370\leq x\leq 2.015)\approx 0.3131$. The model misspecification error for the one-sided test (greater) is almost 12 times higher than that for the one-sided test (less). This means that the one-sided test (less) is more robust than the one-sided test (greater) when $\alpha<0$. Under the settings of $\alpha<0$, one-sided test (less) and not well-mixed observations of single-subject, Crawford et al. (2006) claimed that the $t$ score is robust in terms of Type I error even under skewed Student $t$ distribution. However, this assertion does not hold if the one-sided test (less) is replaced with the one-sided test (greater). For the power comparison, Crawford and Garthwaite (2006) make a similar statement. However, by using a similar discussion to the above, we can find that the robustness statement for the power only holds for the specific settings in their paper.
@@
We can conclude that the interplay between skewness (whether positive or negative), the type of hypothesis (‘two-sided’, ‘less’, or ‘greater’), and the single-subject observations (whether well-mixed or not) significantly influences both the Type I error rate and the power of the $t$ score test under the assumption of the skewed Student $t$ distribution. The key contributions are:
  • This paper introduces a Bayesian Inference General Procedures for A Single-subject Test (BIGPAST) framework, which employs the nested sampling technique under the skewed Student $t$ assumption to fully discuss the test performance in different settings (i.e., the combinations of the sign of $\alpha$, type of hypothesis and well-mixed single-subject observations).
  • We propose a Jeffreys prior, which is defined by explicit mathematical formulas. We evaluate BIGPAST’s performance against the existing approaches through simulation studies. The results demonstrate that BIGPAST is robust against departures from normality and surpasses the existing methods based on the assumption of a Normal distribution in terms of Type I error, power and accuracy.
@@ 19       $i\leftarrow i+1$;
In Algorithm 1, the Metropolis-Hastings algorithm draws samples from the posterior distribution. The step size $\delta$ is set to 0.5, and the burn-in rate $b$ is set to 0.4. The algorithm returns the samples after the burn-in period. $u_{\nu},u_{\omega},u$ are random numbers drawn from the uniform distribution $U(0,1)$. $\epsilon_{\alpha}$ and $\epsilon_{\xi}$ are drawn from the Normal distribution with mean zero and standard deviation $\delta$. A brief explanation of Algorithm 1 is as follows:
  • Lines 6-9: Draw the new parameters $\nu^{\text{new}}$ and $\omega^{\text{new}}$ from the proposal distributions. As $\nu$ and $\omega$ are positive, the proposal distributions for $\nu$ and $\omega$ are the truncated normal distributions with mean $\nu^{\text{old}}$, $\omega^{\text{old}}$ and standard deviation $\delta$, respectively.
  • Lines 10-11: Draw the new parameters $\alpha^{\text{new}}$ and $\xi^{\text{new}}$ from the proposal distributions. The proposal distributions for $\alpha$ and $\xi$ are normal distributions with mean $\alpha^{\text{old}}$, $\xi^{\text{old}}$ and standard deviation $\delta$, respectively.
  • Lines 12-13: Compute the logarithm of the posterior distribution for the new and old parameters.
  • Lines 14-19: Compute the acceptance rate $A$ and update the parameters. The truncated proposal distributions do not have symmetric densities, so we have to adjust the acceptance rate $A$ by adding the logarithm of the ratios, i.e., $p_{\nu}$ and $p_{\omega}$ to the acceptance rate $A$ in line 14.
@@ 2.2 Jeffreys Priors
In the literature, two types of priors are commonly used for $\xi$ and $\omega$ in Equation 2: the partial information prior (Sun and Berger, 1998; Cohen et al., 2025) and the independence Jeffreys prior (Rubio and Steel, 2014; Fortuin, 2022). Both of these priors assume that the joint prior of $\xi$ and $\omega$, denoted as $\boldsymbol{\pi}(\xi,\omega)$, is proportional to $\omega^{-1}$. Consequently, the joint prior of $\alpha,\nu,\xi,\omega$ can be expressed as follows:
@@ 3.3 Model Misspecification Error
In this section, we conduct a comparative analysis of the BIGPAST, CG, and AD methodologies under varying alternative hypotheses and parameter configurations. The predetermined significance level is 0.05. The skewness parameter $\alpha$ and degrees of freedom $\nu$ are assigned values from the sets $(1,10)$, $(1,5)$, $(2,5)$, and $(3,5)$, respectively. The location and scale of the skewed Student $t$ distribution are set to 0 and 1, respectively. Other sample settings include a sample size $n=100$, $c_{\nu}=1$, and a Metropolis-Hastings sampling size of 2000. The comparison procedures are as follows: Firstly, $N=200$ independent samples are randomly generated from the skewed Student $t$ distribution with parameters $\alpha$ and $\nu$. Secondly, $m=1000$ single-subjects are randomly drawn, and the BIGPAST, CG, and AD tests are conducted on each of the $N$ samples. Thirdly, for each of $N$ samples, the Type I error (false positive rate), power (true positive rate), and accuracy are calculated based on the $m=1000$ single-subjects. Finally, the Type I error, power, and accuracy are reported in the form of box plots for the four skewed Student $t$ parameter settings, as shown in Figures 10, 11, 12 and 2.
@@
We utilise the total variation distance to quantify the disparity between skewed Student $t$ (for BIGPAST) and Normal (for CG) distributions. The total variation distances of the aforementioned four settings are: 0.057, 0.111, 0.155 and 0.186, respectively, as illustrated in Figures 10(a), 11(a), 12(a) and 2(a). In the context of a two-sided hypothesis test, the divergence between the BIGPAST and CG methodologies becomes larger as the total variation distance increases, as shown in Figures 10, 11, 12 and 2. The result of the AD approach demonstrates robustness for different total variation distances but performs significantly worse than BIGPAST in terms of Type I error, as seen in Figure 2(b). The AD approach also performs slightly worse than BIGPAST in terms of accuracy, as depicted in Figure 2(d).
@@
The complexity of the one-sided test (either ‘greater’ or ‘less’) compared to the two-sided test due to the skewed Student $t$ distribution increases due to the skewed Student $t$ distribution having light and heavy tails. The model misspecification error impacts the tails of the skewed Student $t$ distribution differently. When $\alpha>0$ (resp. $\alpha<0$), the skewed Student $t$ distribution has a right (resp. left) heavy tail. In this paragraph, we explore the scenario where the direction of the alternative hypothesis is ‘greater’, i.e., one-sided test (greater). If we implement the CG test, the 95% quantile of the Normal distribution is less than the 95% quantile of the skewed Student $t$ distribution. Consequently, the Type I error of the CG test exceeds that of the BIGPAST test, leading to a higher power for the CG test, as shown in Figures 2(b) and 2(c). This can be explained by the light-red region between the 95% quantile of skewed Student $t$ distribution and the 95% quantile of Normal distribution, as depicted in Figure 3(b). Theoretically, all single-case observations from this region will be classified as positive by the CG test and as negative by the BIGPAST test, leading us to refer to this region as the false positive region. The error generated by the false positive region arises from model misspecification rather than sampling randomness, as illustrated in Figure 3.
@@ 4 Real Data Analysis
This dataset is derived from a mild traumatic brain injury (mTBI) study conducted by Innovision IP Ltd. It comprises a functional time series of Magnetoencephalography (MEG) data and an anatomical T1-weighted Magnetic Resonance Imaging (MRI) scan for 103 healthy subjects and one patient with mild traumatic brain injury. The MEG data, recorded from 306 sensors, was pre-processed using the ‘MNE-Python’ package (Gramfort et al., 2013). The MEG data were analysed to reconstruct the source generators using a minimum norm inverse method. The MRI data were pre-processed with the ‘FreeSurfer’ package (Han et al., 2006). The standard FreeSurfer pipeline (detailed on their website) was followed and includes motion correction, intensity normalisation, skull-stripping, normalisation to template space and segmentation of the cortex and subcortical structures using the Desikan-Killiany atlas. This results in 68 regions covering the entire cortex. For each cortical region, we calculate the average of source powers within this region. This process yields a control group dataset with dimensions $103\times 68\times 100$ and a patient dataset with dimensions $68\times 100$. To simplify, we treat the epochs as repeated observations and take averages over them to reduce the data’s complexity. Consequently, the control group dataset has dimensions $103\times 68$, and the patient dataset is reduced to a vector of length $68$.